# Zero-Shot Scene Change Detection

### v1

@@ Abstract
We present a novel, training-free approach to scene change detection. Our method leverages tracking models, which inherently perform change detection between consecutive frames of video by identifying common objects and detecting new or missing objects. Specifically, our method takes advantage of the change detection effect of the tracking model by inputting reference and query images instead of consecutive frames. Furthermore, we focus on the content gap and style gap between two input images in change detection, and address both issues by proposing adaptive content threshold and style bridging layers, respectively. Finally, we extend our approach to video to exploit rich temporal information, enhancing scene change detection performance. We compare our approach and baseline through various experiments. While existing train-based baseline tend to specialize only in the trained domain, our method shows consistent performance across various domains, proving the competitiveness of our approach.
Keywords: Scene Change Detection, Zero-shot learning
1 Introduction
Scene Change Detection (SCD) is the task that aims to detect differences between two scenes separated by a temporal interval, and it is a fundamental capability required for mobile robots to ensure reliable navigation, and task performance completion. If a robot lacks the capability of SCD, several problems could arise, affecting its performance, safety, and overall effectiveness. Without SCD, for instance, a robot might not be able to detect recently added barriers or changes in the environment, which could result in collisions and possible harm to the robot as well as other objects. Furthermore, without the capability of SCD, the robot’s maps of the environment would not be updated frequently enough, reducing the accuracy of localization and navigation tasks. This can lead to increased errors in estimating its position. Conversely, robots equipped with SCD capabilities can be employed in various applications, such as disaster detection [1, 2], terrain monitoring [3, 4, 5], and industrial warehouse management [6, 7].
@@ 1 Introduction
Recently, SCD has been tackled using deep learning. Deep learning-based SCD techniques follow a procedure of learning from a training dataset and applying the model to a test dataset. These approaches tend to face two main challenges: dataset generation costs and susceptibility to style variations. Firstly, creating a training dataset for SCD models is labor-intensive and costly. Recent research has focused on reducing these costs through semi-supervised [8, 9] and self-supervised learning [10, 11] methods, as well as the use of synthetic data [12, 8]. While these approaches mitigate the expense of labeling, they often overlook the cost of acquiring image pairs and fail to address the unnaturalness of collecting such data, as humans typically do not need image pairs for change detection. Secondly, due to the substantial temporal intervals between pre-change and post-change images, variations in seasons, weather, and time introduce significant differences in their visual characteristics. Consequently, SCD techniques must be robust to these style variations to be effective. However, the training dataset cannot include all the style variations present in real-world scenarios, making the trained model vulnerable to style variations that are not included in the training set.
To address these problems, we propose a novel training-free zero-shot SCD method. Our method does not require a training dataset, thereby eliminating collection costs and allows it to be applied to any problem with arbitrary styles. To the best of our knowledge, this paper is the first to attempt zero-shot SCD without training on a training dataset. The key idea of this paper is to formulate SCD as a tracking problem, and apply a foundation tracking model to conduct zero-shot SCD. This idea stems from the observation that the tracking task is fundamentally similar to change detection. Specifically, tracking models [13, 14] maintain or build tracks by identifying the same objects, disappeared objects, and newly appeared objects in two consecutive images, even when the camera and objects move. Thus, if the two consecutive images in tracking are replaced with the two images before and after the change in SCD, the tracking model can automatically solve SCD without training.
However, it is worth noting that there are some differences between tracking and SCD: (a) Unlike tracking, the two images before and after the change in SCD might have different styles due to a large time gap between two images. We refer to this SCD trait as the style gap. (b) Objects change very little between two consecutive images in tracking, whereas objects change abruptly in SCD. We refer to this SCD trait as the content gap. To address these issues in our zero-shot SCD method, we introduce a style bridging layer and a content threshold, respectively.
Finally, we extend our approach to sequence, to introduce the zero-shot SCD approach works on the video. Since our approach operates based on a tracking model, it can be seamlessly extended to work with video sequences. The proposed zero-shot SCD approach has been evaluated on three benchmark datasets and shows that our zero-shot SCD demonstrates comparable or even superior performance compared to previous state-of-the-art training-based SCD methods.
2 Related Work
Scene Change Detection (SCD): In recent years, numerous deep learning-based change detection methods have been proposed for SCD. DR-TANet [15] utilized attention mechanism based on the encoder-decoder architecture. SimSaC [7] developed a network with a warping module to correct distortions between images. C-3PO [16] developed a network that fuses temporal features to distinguish three change types. However, various studies have aimed to address the challenge of obtaining data. For instance, [8, 9] have introduced semi-supervised learning, and [10, 11] proposed the self-supervised learning with unlabeled data. However, [12, 8] utilized synthetic data to effectively increase the dataset. Despite these methods effectively reducing the label costs, they tend to overlook the cost of collecting image pairs. Moreover, the robustness against style change has not been previously discussed. An effective SCD method should be able to focus on content changes regardless of variations in image style. However, since a single dataset cannot encompass all possible style variations, the performance tends to be specialized for the styles present in the dataset. This issue, while not evident in controlled laboratory environments, becomes a significant problem in real-world applications. Therefore, we propose a novel SCD method that does not rely on datasets. Our method operates without a training dataset, thus ensuring independence from specific styles.
Segmenting and Tracking Anything: Recently, Segment Anything [17] has demonstrated highly effective performance in universal image segmentation. SAM has shown the ability to perform various zero-shot tasks, and has served as the foundational model for various studies [18, 19]. Building upon this research, researchers have explored various methods to extend its application to tracking. For example, SAM-Track [20] implemented tracking by combining SAM with the DeAOT [21] mask tracker. SAM-PT [22] integrated SAM with point tracking to develop the pipeline. DEVA [13] proposed a pipeline that uses the XMem tracker [14] to track provided masks without additional training. Among various studies, we adopted DEVA with SAM masks as our tracking model, to achieve track-anything for SCD without further training.
@@ 2 Related Work
Figure 1: The Basic idea of SCD with tracking model. (a) We execute the tracking model $G$ with $r$ and $q$. (b) We denote the tracking result from $r$ to $q$ as $M^{r\to q}=G(r,q,M^{r})$, and the tracking result from $q$ to $r$ as $M^{q\to r}=G(q,r,M^{q})$ (c) ‘Missing’ objects are the objects that exist in $r$ but not in $q$. Therefore, we compare $M^{r}$ and $M^{r\to q}$ to find missing objects. Conversely, ‘new’ objects are identified by comparing $M^{q}$ and $M^{q\to r}$. (d) The final prediction is the simple combination of new and missing.
3 Method
@@
To perform scene change detection without training, our methodology integrates two pretrained models: a segmentation model $F$ and a tracking model $G$. The segmentation model $F$ segments images in an unsupervised manner, while the tracking model $G$ tracks each mask generated by $F$ across multiple images. We employ the Segment Anything Model (SAM) [17] as the segmentation model $F$ and DEVA [13] for the tracking model $G$. Comprehensive details about parameters for $F$ and $G$, and details about mask generation process will be provided in the supplementary materials.
We first introduce the basic idea for performing SCD between two images using $F$ and $G$ in Section 3.1. Next, we discuss the differences between the tracking task and the SCD task, and then introduce methods to overcome these differences in Section 3.2. Finally, we extend our approach to the video level in Section 3.3.
3.1 Scene Change Detection with Tracking model
Our approach uses two pretrained models, a segmentation model $F$ and a tracking model $G$. The segmentation model $F$ partitions image $I$ into object-level masks, forming the set $M=\{m_{0},m_{1},\cdots,m_{n}\}$. There exists no overlap between distinct masks, that is, $m_{i}\cap m_{j}=0,\forall~{}i\neq j$. The tracking model $G$ takes consecutive frame images $(I^{0},I^{1})$ and the object masks of the first frame $M^{0}=F(I^{0})$ as input, and yields $M^{0\to 1}$ as output, that is, $M^{0\to 1}=G(I^{0},I^{1},M^{0})$. Here, $M^{0\to 1}$ represents the set of masks tracked from $M^{0}=F(I^{0})$ to $I^{1}$. By checking if each mask that was present in $M^{0}$ is also present in $M^{0\to 1}$, we can see which masks still exist and which ones have disappeared from $I^{0}$ and $I^{1}$.
The key idea of our zero-shot SCD is to apply a reference frame $r$ and a query frame $q$ instead of consecutive frames $(I^{0},I^{1})$ to the tracking model $G$. Although the tracking model traditionally expects consecutive frames $(I^{0},I^{1})$ for input, we deviate from this convention by providing reference frame $r$ and query frame $q$ instead. To avoid potential confusion, we rewrite the input and the output of the tracking model as $M^{r\to q}=G(r,q,M^{r})$. By comparing the masks between $M^{r}$ and $M^{r\to q}$, we identify object masks that exist at time $t_{0}$ but have disappeared at time $t_{1}$, corresponding to the ‘missing’ class in the change detection task. Specifically,
@@ 3.1 Scene Change Detection with Tracking model
Additionally, we run the tracking model $G$ again by reversing the order reference frame $r$ and query frame $q$ and feed them to $G$ to obtain $M^{q\to r}=G(q,r,M^{q})$. Similarly, we predict the ‘new’ objects by $M_{new}=M^{q}\setminus M^{q\to r}$, which represent the objects that appear at time $t_{1}$ but were absent at time $t_{0}$. Our pixel-wise prediction is obtained by applying the union of masks within $M_{new}$ and $M_{missing}$. Pixels experiencing both ‘new’ and ‘missing’ occurrences are considered ‘replaced.’ Formally, change prediction $P_{changed}$ is determined by:
@@
3.2 Addressing Content Gap and Style Gap
As presented in Section 3.1, the key idea of our zero-shot SCD is to exploit the similarity between SCD and tracking tasks. However, directly applying this concept to various SCD scenarios can sometimes result in performance degradation due to inherent differences between the two tasks. In this section, we examine these differences and propose corresponding solutions.
@@ 3.2 Addressing Content Gap and Style Gap
The first difference is the content gap, which refers to abrupt changes in content between the reference and query images. In traditional tracking tasks, objects typically disappear gradually over multiple frames rather than suddenly, and new objects appear gradually over multiple frames, implying that tracking tasks have little content gap. In contrast, SCD involves abrupt changes where objects disappear or appear within a single frame and it has a large content gap. Therefore, when the tracking model $G$ trained on the tracking dataset is directly applied to SCD, the tracking model $G$ tends to create small segments even for objects that have disappeared, as shown in Figure 2. In the first row, the yellow forklift in the query image is missing from the reference image, but, in the second row, the mask in $M^{q\to r}$ tracked from a blue mask in $M^{q}$ has a small segment. This remaining small segment from the tracking model $G$ makes the identification of missing objects very difficult.
To address the problem, when the size of an object is significantly reduced after the tracking, we consider the object has disappeared even if its size is not completely vanished. To the end, we introduce a content threshold $\tau$ and we compare the areas of the masks before and after tracking. If the ratio is less than the content threshold $\tau$, we consider the corresponding object is missing or newly appeared. We define the $\setminus\!^{\tau}$ operator to replace the $\setminus$ operator in equation 1 as follows, where $|m^{*}_{i}|$ denotes the area of the $i$-th mask in set *:
$A\setminus\!\!^{\tau}\,B:=\{m^{A}_{i}\,|\,{\,|m^{B}_{i}|\,\over\,|m^{A}_{i}|\,}<\tau,\quad\forall m^{A}_{i}\in A\}.$ (3)
By introducing this operator, we have $M_{missing}=M^{r}\setminus\!^{\tau}\,M^{r\to q}$. The value of $\tau$ is set to 0.05.
The second difference between tracking tasks and SCD is the style gap, which refers to the difference in style between the reference and query images. This style gap arises in SCD because there is a long time interval between the capture of the reference image and the query image, during which weather, time of day, or even season can change. Such variations are not considered in traditional tracking tasks and it can significantly degrade the SCD performance.
@@
To address the style gap, we introduce a style bridging layer (SBL) by incorporating an Adaptive Instance Normalization (AdaIN) layer [23] into the residual blocks of ResNet backbone [24]. The AdaIN layer, widely used to reduce style differences between images across various fields [25, 26, 27], references the first image and applies its style to the second image, thereby reducing the style differences between the two images. Inspired by this, SBL addresses the style gap between two inputs without learning, with two training-free style parameters. For example, during the process $M^{r\to q}=G(r,q,M^{r})$, the style bridging layer records the mean and variance of each layer in image $r$ and applies these statistics when processing image $q$. Formally, the style bridging layer updates the feature of $z^{q}_{l}$ as follows, where $z^{*}_{l}$ denotes the $l$-th layer feature of image *.
$\tilde{z}^{q}_{l}=\sigma(z^{r}_{l}){{z^{q}_{l}-\mu(z^{q}_{l})}\over{\sigma(z^{q}_{l})}}+\mu(z^{r}_{l}).$ (4)
@@
3.3 Extension to the Video Sequences
In this subsection, the SCD between two images is extended to the SCD between two videos. Using the tracking model enables us to seamlessly extend image SCD to video SCD. Consider a video SCD dataset consisting of sequences of reference, query, and change labels, denoted as $\{r^{t},q^{t},y^{t}\}_{t=1}^{T}$, where $T$ represents the length of the video sequence, and $t$ denotes the time index. Utilizing the videos in SCD provides a richer representation of the objects and improves the SCD performance. Compared to the image SCD, the video SCD requires three modifications. The first modification is simply to feed two sequences $\{r^{t},q^{t}\}_{t=1}^{T}$ instead of image pair $\{r,q\}$ as input to the tracking model. Specifically, we start the video SCD with
@@ 3.3 Extension to the Video Sequences
and proceed with our SCD by
@@
where $M^{r^{1}\twoheadrightarrow r^{t}}=M^{r^{1}\to r^{2}\to\cdots\to r^{t-1}\to r^{t}}$ and $M^{r^{1}\twoheadrightarrow q^{t}}=M^{r^{1}\to r^{2}\to\cdots\to r^{t-1}\to r^{t}\to q^{t}}$, as shown in Figure 4. In mask formulation, $M^{X}$ denotes the output from the segmentation model $F$, whereas $M^{X\to Y}$ or $M^{X\twoheadrightarrow Y}$ denotes the output from tracking model $G$. This architecture is slightly similar to the structure of the Bayes filters or Markov process [28], in that all the information processed from $1$ to $t-1$ are included in $M^{r^{1}\twoheadrightarrow r^{t-1}}$. Consequently, $M^{r^{1}\twoheadrightarrow r^{t}}$ can be incrementally updated from $M^{r^{1}\twoheadrightarrow r^{t-1}}$ and ${r^{t},q^{t}}$, without reprocessing all previous images. During the incremental update from $M^{r^{1}\twoheadrightarrow r^{t-1}}$ to $M^{r^{1}\twoheadrightarrow r^{t}}$, all functions for tracking, including updating feature memory and discovering new objects, are activated. Conversely, during the update from $M^{r^{1}\twoheadrightarrow r^{t}}$ to $M^{r^{1}\twoheadrightarrow q^{t}}$, these functions are deactivated. This processing sequence is designed to detect missing objects, whereas the opposite processing sequence with swapping $r$ and $q$ is employed to detect new objects.
@@
The second modification is to define $M^{t}_{missing}$ and $M^{t}_{new}$ at time $t$. To find missing objects, we consider all frames in query sequence $\{q^{t}\}_{t=1}^{T}$, as follows:
$M_{missing}^{t}=\{M^{r^{t}}\setminus\!\!^{\tau}\,M^{r^{1}\twoheadrightarrow q^{1}}\}\cap\{M^{r^{t}}\setminus\!\!^{\tau}\,M^{r^{1}\twoheadrightarrow q^{2}}\}\cap\cdots\cap\{M^{r^{t}}\setminus\!\!^{\tau}\,M^{r^{1}\twoheadrightarrow q^{T}}\}.$ (7)
This definition of $M^{t}_{missing}$ indicates that the missing object is the object that exists in reference frame $r^{t}$, but is absent in the query sequence $\{q^{t}\}_{t=1}^{T}$. $M^{t}_{new}$ is similarly defined with the object that exists in query frame $q^{t}$, but is absent in the reference sequence $\{r^{t}\}_{t=1}^{T}$.
However, with more cross-sequence comparisons, there is a higher noise exposure. Therefore, as the sequence length increases, it becomes necessary to adjust the content threshold introduced in Section 3.2. The third modification is extending it to an adaptive content threshold based on the sequence length, defined by the following equation:
$\tau=0.5-{{0.9}\over{\sqrt{length}+1}}.$ (8)
As the sequence length increases, the threshold becomes higher. The threshold is capped at 0.5.
Through these three extensions, our methodology becomes appropriate for processing videos. According to the definitions of these three modifications, the sequence length can range from 1 to infinity. However, we impose an upper bound on the length of a sequence, denoted as $T_{max}$. If the length of the video exceeds $T_{max}$, it is treated as multiple sequences each with a length of $T_{max}$ individually. The reason for constraining the length of sequences is simple: as sequences lengthen, memory costs increase, while the relevant information for change detection diminishes. For instance, in scenarios where the camera is in motion, the initial and final frames of a sequence may capture entirely different locations, rendering them unsuitable for change detection. Conversely, if the camera remains stationary, all frames depict the same scene, and redundant frames do not contribute useful information. Therefore, we ensure more effective SCD by setting the upper bound of the sequence length. For our experiments, we set $T_{max}$ to 60.
4 Experiments
4.1 Experimental Setup
In this section, we briefly introduce the dataset, the relevant settings, and the evaluation metrics.
ChangeSim [6] is a synthetic dataset with an industrial indoor environment. It includes three subsets with varying environmental conditions: normal, low-illumination, and dusty air. The dataset categorizes changes into four types: new, missing, rotated, and replaced. Despite its variety of environmental variations and change classes, most baseline experiments on this dataset have evaluated only the binary change/unchange classification and have predominantly focused on the normal subset, leaving the dataset’s full potential underexplored. Therefore, we chose the state-of-the-art method, C-3PO [16], and reproduced the results under the following conditions: using the original image size ($640\times 480$) to fully utilize the rich information; and including all three subsets. Among the four change classes in this dataset, the rotated class, unlike others, involves slight angular changes of the same object rather than complete appearances or disappearances. We considered this as the object remaining static and integrated this class into static for evaluation.
Evaluation Metrics. We employ the mean Intersection over Union (mIoU) metric.
4.2 Experimental Results

### v2

@@ Abstract
We present a novel, training-free approach to scene change detection. Our method leverages tracking models, which inherently perform change detection between consecutive frames of video by identifying common objects and detecting new or missing objects. Specifically, our method takes advantage of the change detection effect of the tracking model by inputting reference and query images instead of consecutive frames. Furthermore, we focus on the content gap and style gap between two input images in change detection, and address both issues by proposing adaptive content threshold and style bridging layers, respectively. Finally, we extend our approach to video, leveraging rich temporal information to enhance the performance of scene change detection. We compare our approach and baseline through various experiments. While existing train-based baseline tend to specialize only in the trained domain, our method shows consistent performance across various domains, proving the competitiveness of our approach.
Codehttps://github.com/kyusik-cho/ZSSCD
Introduction
Scene Change Detection (SCD) is the task that aims to detect differences between two scenes separated by a temporal interval. Recently, SCD has gained significant interest in various applications involving mobile drones and robots. For instance, drone-based SCD has been studied for land terrain monitoring (Lv et al. 2023; Song et al. 2019; Agarwal, Kumar, and Singh 2019), construction progress monitoring (Han et al. 2021), and urban feature monitoring (Chen et al. 2016). Additionally, SCD using mobile robots has also been researched for natural disaster damage assessment (Sakurada and Okatani 2015; Sakurada, Okatani, and Deguchi 2013), urban landscape monitoring (Alcantarilla et al. 2018; Sakurada, Shibuya, and Wang 2020), and industrial warehouse management (Park et al. 2021, 2022).
@@ Introduction
Recently, SCD has been tackled using deep learning. Deep learning-based SCD techniques follow a procedure of learning from a training dataset and applying the model to a test dataset. These approaches tend to face two main challenges: dataset generation costs and susceptibility to style variations. Firstly, creating a training dataset for SCD models is labor-intensive and costly. Recent research has focused on reducing these costs through semi-supervised (Lee and Kim 2024; Sun et al. 2022) and self-supervised learning (Seo et al. 2023; Furukawa et al. 2020) methods, as well as the use of synthetic data (Sachdeva and Zisserman 2023; Lee and Kim 2024). While these approaches mitigate the expense of labeling, they often overlook the cost of acquiring image pairs, which arises from the substantial temporal intervals to capture changes. Secondly, due to the substantial temporal intervals between pre-change and post-change images, variations in seasons, weather, and time introduce significant differences in their visual characteristics. Consequently, SCD techniques must be robust to these style variations to be effective. However, the training dataset cannot include all the style variations present in real-world scenarios, making the trained model vulnerable to style variations that are not included in the training set.
@@
Figure 1: The Basic idea of SCD with tracking model. (a) We execute the tracking model $G$ with $r$ and $q$. (b) We denote the tracking result from $r$ to $q$ as $M^{r\to q}=G(r,q,M^{r})$, and the tracking result from $q$ to $r$ as $M^{q\to r}=G(q,r,M^{q})$. (c) ‘Missing’ objects are the objects that exist in $r$ but not in $q$. Therefore, we compare $M^{r}$ and $M^{r\to q}$ to find missing objects. Conversely, ‘new’ objects are identified by comparing $M^{q}$ and $M^{q\to r}$. (d) The final prediction is the simple combination of new and missing.
To address these problems, we propose a novel training-free zero-shot SCD method. Our method does not require a training dataset, thereby eliminating collection costs and allows it to be applied to any problem with arbitrary styles. To the best of our knowledge, this paper is the first to attempt zero-shot SCD without training on SCD datasets. The key idea of this paper is to formulate SCD as a tracking problem, and apply a foundation tracking model to conduct zero-shot SCD. This idea stems from the observation that the tracking task is fundamentally similar to change detection. Specifically, tracking models (Cheng et al. 2023a; Cheng and Schwing 2022) maintain or build tracks by identifying the same objects, disappeared objects, and newly appeared objects in two consecutive images, even when the camera and objects move. Thus, if the two consecutive images in tracking are replaced with the two images before and after the change in SCD, the tracking model can automatically solve SCD without training.
However, there are some differences between tracking and SCD tasks: (a) Unlike tracking, the two images before and after the change in SCD might have different styles due to a large time gap between two images. We refer to this SCD trait as the style gap. (b) Objects change very little between two consecutive images in tracking, whereas objects change abruptly in SCD. We refer to this SCD trait as the content gap. To address these issues in our zero-shot SCD method, we introduce a style bridging layer and a content threshold, respectively.
Finally, we extend our approach to sequence, to introduce the zero-shot SCD approach works on the video. Since our approach operates based on a tracking model, it can be seamlessly extended to work with video sequences. The proposed zero-shot SCD approach has been evaluated on three benchmark datasets and demonstrated comparable or even superior performance compared to previous state-of-the-art training-based SCD methods.
Related Work
Scene Change Detection (SCD)
In recent years, numerous deep learning-based change detection methods have been proposed for SCD. DR-TANet (Chen, Yang, and Stiefelhagen 2021) utilized attention mechanism based on the encoder-decoder architecture. SimSaC (Park et al. 2022) developed a network with a warping module to correct distortions between images. C-3PO (Wang, Gao, and Wang 2023) developed a network that fuses temporal features to distinguish three change types. However, various studies have aimed to address the challenge of obtaining data. For instance, Lee and Kim (2024) and Sun et al. (2022) have introduced semi-supervised learning, while Seo et al. (2023) and Fukawa et al. (2020) proposed the self-supervised learning with unlabeled data. However, Sachdeva and Zisserman (2023) and Lee and Kim (2024) utilized synthetic data to effectively increase the dataset. Despite these methods effectively reducing the label costs, they tend to overlook the cost of collecting image pairs. Moreover, the robustness against style change has not been previously discussed. An effective SCD method should be able to focus on content changes regardless of variations in image style. However, since a single dataset cannot encompass all possible style variations, the performance tends to be specialized for the styles present in the dataset. This issue, while not evident in controlled laboratory environments, becomes a significant problem in real-world applications. Therefore, we propose a novel SCD method that does not rely on datasets. Our method operates without a training dataset, thus ensuring independence from specific styles.
Segmenting and Tracking Anything
Recently, Segment Anything (Kirillov et al. 2023) has demonstrated highly effective performance in universal image segmentation. SAM has shown the ability to perform various zero-shot tasks, and has served as the foundational model for various studies (Peng et al. 2023; Maquiling et al. 2024). Building upon this research, researchers have explored various methods to extend its application to tracking. For example, SAM-Track (Cheng et al. 2023b) implemented tracking by combining SAM with the DeAOT (Yang and Yang 2022) mask tracker. SAM-PT (Rajič et al. 2023) integrated SAM with point tracking to develop the pipeline. DEVA (Cheng et al. 2023a) proposed a pipeline that uses the XMem tracker (Cheng and Schwing 2022) to track provided masks without additional training. Among various studies, we adopted DEVA with SAM masks as our tracking model, to achieve track-anything for SCD without further training.
Method
@@
To perform scene change detection without training, our methodology integrates two pretrained models: a segmentation model $F$ and a tracking model $G$. The segmentation model $F$ segments images in an unsupervised manner, while the tracking model $G$ tracks each mask generated by $F$ across multiple images. We employ the Segment Anything Model (SAM) (Kirillov et al. 2023) as the segmentation model $F$ and DEVA (Cheng et al. 2023a) for the tracking model $G$. Comprehensive details about parameters for $F$ and $G$, and details about the mask generation process are provided in the supplementary materials.
The rest of the Method section is structured as follows: First, we introduce the basic idea for performing SCD between two images using $F$ and $G$. Next, we discuss the differences between the tracking task and the SCD task, and then introduce methods to overcome these differences. Finally, we extend our approach to the video level.
Scene Change Detection with Tracking model
Our approach uses two pretrained models, a segmentation model $F$ and a tracking model $G$. The segmentation model $F$ partitions image $I$ into object-level masks, forming the set $M=\{m_{1},m_{2},\cdots,m_{n}\}$. There exists no overlap between distinct masks, that is, $m_{i}\cap m_{j}=0,\forall~{}i\neq j$. The tracking model $G$ takes consecutive frame images $(I^{0},I^{1})$ and the object masks of the first frame $M^{0}=F(I^{0})$ as input, and yields $M^{0\to 1}$ as output, that is, $M^{0\to 1}=G(I^{0},I^{1},M^{0})$. Here, $M^{0\to 1}$ represents the set of masks tracked from $M^{0}=F(I^{0})$ to $I^{1}$. By checking if each mask that was present in $M^{0}$ is also present in $M^{0\to 1}$, we can determine which object masks in $I^{0}$ still exist or have disappeared in $I^{1}$.
The key idea of our zero-shot SCD approach is to apply a reference image $r$ and a query image $q$ instead of consecutive frames $(I^{0},I^{1})$ to the tracking model $G$. Although the tracking model traditionally expects consecutive frames $(I^{0},I^{1})$ for input, we deviate from this convention by providing reference image $r$ and query image $q$ instead. To avoid potential confusion, we rewrite the input and the output of the tracking model as $M^{r\to q}=G(r,q,M^{r})$. By comparing the masks between $M^{r}$ and $M^{r\to q}$, we identify object masks that exist at time $t_{0}$ but have disappeared at time $t_{1}$, corresponding to the ‘missing’ class in the change detection task. Specifically,
@@
Additionally, we run the tracking model $G$ again by reversing the order reference image $r$ and query image $q$ and feed them to $G$ to obtain $M^{q\to r}=G(q,r,M^{q})$. Similarly, we predict the ‘new’ objects by $M_{new}=M^{q}\setminus M^{q\to r}$, which represent the objects that appear at time $t_{1}$ but were absent at time $t_{0}$. Our pixel-wise prediction is obtained by applying the union of masks within $M_{new}$ and $M_{missing}$. Pixels experiencing both ‘new’ and ‘missing’ occurrences are considered ‘replaced.’ Formally, change prediction $P_{changed}$ is determined by:
@@
Addressing Content Gap and Style Gap
As presented in the previous section, the key idea of our zero-shot SCD approach is to exploit the similarity between SCD and tracking tasks. However, directly applying this concept to various SCD scenarios leads to suboptimal performance due to inherent differences between the two tasks. In this section, we analyze these differences and propose corresponding solutions.
The first difference is the content gap, which refers to abrupt changes in content between the reference and query images. In traditional tracking tasks, objects typically disappear gradually over multiple frames rather than suddenly, and new objects appear gradually over multiple frames, implying that tracking tasks have little content gap. In contrast, SCD involves abrupt changes where objects disappear or appear within a single frame and it has a large content gap. Therefore, when the tracking model $G$ trained on the tracking dataset is directly applied to SCD, the tracking model $G$ tends to create small segments even for objects that have disappeared, as shown in Figure 2. In the first row, the yellow forklift in the query image is missing from the reference image, but, in the second row, the mask in $M^{q\to r}$ tracked from a blue mask in $M^{q}$ has a small segment. This remaining small segment makes the identification of missing objects very difficult.
To address the content gap, we propose considering an object as disappeared if its size is significantly reduced after tracking, even if it has not completely vanished. To the end, we introduce a content threshold $\tau$ and we compare the areas of the masks before and after tracking. If the ratio is less than the content threshold $\tau$, we consider the corresponding object is missing or newly appeared. We define the $\setminus\!^{\tau}$ operator to replace the $\setminus$ operator in equation 1 as follows, where $|m^{*}_{i}|$ denotes the area of the $i$-th mask in set *:
$A\setminus\!\!^{\tau}\,B:=\{m^{A}_{i}\,|\,{\frac{|m^{B}_{i}|}{\,|m^{A}_{i}|\,}}<\tau,\quad\forall m^{A}_{i}\in A\}.$ (3)
By introducing this operator, we have $M_{missing}=M^{r}\setminus\!^{\tau}\,M^{r\to q}$. The value of $\tau$ is automatically determined based on the input data. Further details and discussions on the determination of $\tau$ are provided in the supplementary material.
The second difference between tracking tasks and SCD is the style gap, which refers to the difference in style between the reference and query images. Change detection data have large temporal gaps, therefore lighting, weather, or season can change. These changes are commonly modeled as style changes (Tang et al. 2023). We define such variations as the style gap, which is not considered in traditional tracking tasks and can thus significantly degrade SCD performance.
@@
To address the style gap, we introduce a style bridging layer (SBL) by incorporating an Adaptive Instance Normalization (AdaIN) layer (Huang and Belongie 2017) into the residual blocks of ResNet backbone (He et al. 2016) of the tracking model $G$. The AdaIN layer, widely used to reduce style differences between images across various fields (Karras, Laine, and Aila 2019; Wu et al. 2019; Xu et al. 2023), references the first image and applies its style to the second image, thereby reducing the style differences between the two images. Inspired by this, SBL addresses the style gap between two inputs without learning, with two training-free style parameters. For example, during the process $M^{r\to q}=G(r,q,M^{r})$, the style bridging layer records the mean and variance of each layer in image $r$ and applies these statistics when processing image $q$. Formally, the style bridging layer updates the feature of $z^{q}_{l}$ as follows, where $z^{*}_{l}$ denotes the $l$-th layer feature of image *.
$\tilde{z}^{q}_{l}=\sigma(z^{r}_{l}){\frac{\,z^{q}_{l}-\mu(z^{q}_{l})\,}{\sigma(z^{q}_{l})}}+\mu(z^{r}_{l}).$ (4)
@@
Extension to the Video Sequences
In this subsection, we extend our image-based SCD approach to video sequences. Leveraging video data enhances spatial understanding by capturing scenes from multiple angles. The tracking model in our pipeline allows the seamless extension of our image-based SCD approach to video data.
Consider a video SCD dataset consisting of sequences of reference, query, and change labels, denoted as $\{r^{t},q^{t},y^{t}\}_{t=1}^{T}$, where $T$ represents the length of the video sequence, and $t$ denotes the time index. Compared to the image SCD, the video SCD requires two modifications. The first modification is simply to feed two sequences $\{r^{t},q^{t}\}_{t=1}^{T}$ instead of image pair $\{r,q\}$ as input to the tracking model. Specifically, we start tracking to detect missing objects in the video with
@@
and continue tracking throughout the entire video with
@@
where $M^{r^{1}\twoheadrightarrow r^{t}}=M^{r^{1}\to r^{2}\to\cdots\to r^{t-1}\to r^{t}}$ and $M^{r^{1}\twoheadrightarrow q^{t}}=M^{r^{1}\to r^{2}\to\cdots\to r^{t-1}\to r^{t}\to q^{t}}$, as shown in Figure 4. In mask formulation, $M^{X}$ denotes the output from the segmentation model $F$, whereas $M^{X\to Y}$ or $M^{X\twoheadrightarrow Y}$ denotes the output from tracking model $G$. This architecture is similar to the structure of the Bayes filters or Markov process (Thrun 2002), in that all the information processed from $1$ to $t-1$ are included in $M^{r^{1}\twoheadrightarrow r^{t-1}}$. Consequently, $M^{r^{1}\twoheadrightarrow r^{t}}$ can be incrementally updated from $M^{r^{1}\twoheadrightarrow r^{t-1}}$ and ${r^{t},q^{t}}$, without reprocessing all previous images. During the incremental update from $M^{r^{1}\twoheadrightarrow r^{t-1}}$ to $M^{r^{1}\twoheadrightarrow r^{t}}$, all functions for tracking, including updating feature memory and identifying new objects, are activated. Conversely, during the update from $M^{r^{1}\twoheadrightarrow r^{t}}$ to $M^{r^{1}\twoheadrightarrow q^{t}}$, these functions are deactivated. This processing sequence is designed to detect ‘missing’ objects, whereas the opposite processing sequence with swapping $r$ and $q$ is employed to detect ‘new’ objects.
@@
The second modification is redefining $M^{t}_{missing}$ and $M^{t}_{new}$ to suit video data. In equation 1, we defined $M_{missing}$ in image SCD as the masks that exist in the reference image $r$ but are absent in the query image $q$. The definition of $M^{t}_{missing}$ for video SCD is an extension of this definition to video. Specifically, $M^{t}_{missing}$ is defined as the masks present in the reference sequence $\{r^{t}\}_{t=1}^{T}$ but absent in the query sequence $\{q^{t}\}_{t=1}^{T}$. Formally, $M^{t}_{missing}$ at time index $t$ is defined by:
$M_{missing}^{t}=$$\;\;\{M^{r^{t}}\setminus\!\!^{\tau}\,M^{r^{1}\twoheadrightarrow q^{1}}\}$ (7)
$\cap\{M^{r^{t}}\setminus\!\!^{\tau}\,M^{r^{1}\twoheadrightarrow q^{2}}\}$
[…]
$\cap\{M^{r^{t}}\setminus\!\!^{\tau}\,M^{r^{1}\twoheadrightarrow q^{T}}\}.$
$M^{t}_{new}$ is similarly defined with the object that exists in query sequence $\{q^{t}\}_{t=1}^{T}$, but is absent in the reference sequence $\{r^{t}\}_{t=1}^{T}$.
Through these two extensions, our methodology becomes appropriate for processing videos. According to the definitions of these modifications, the sequence length can range from 1 to infinity. However, we impose an upper bound on the length of a sequence, denoted as $T_{max}$. If the length of the video exceeds $T_{max}$, it is divided into multiple sequences, each with a length of $T_{max}$. The reason for constraining the length of sequences is simple: as sequences lengthen, memory costs increase, while the relevant information for change detection decreases. For instance, in scenarios where the camera is in motion, the initial and final frames of a sequence may capture entirely different locations, rendering them unsuitable for change detection. Conversely, if the camera remains stationary, all frames depict the same scene, and additional frames provide redundant information. Therefore, we ensure more effective SCD by setting the upper bound of the sequence length. For our experiments, we set $T_{max}$ to 60.